An investment firm offers its customers municipal bonds that mature after varying numbers of years. Given that the cumulative distribution function of Y, the number of years to maturity for a randomly selected bond, is 4/5, 5 1,y у< 7; find (a) P(Y 5) (b) P(Y > 3) (d) P(Y-51 Y-2)
Answer #1

a)P(Y=5)=P(Y<=5)-P(Y=5)=4/5-3/5=1/5

b)P(Y>3) =1-P(Y<=3)=1-3/5=2/5

c)P(1.4<Y<6)=P(Y<=5)-P(Y<3)=4/5-1/5=3/5

d)P(Y<=5|Y>=2)=P(2<=Y<=5)/P(Y>=2)=(4/5-1/5)/(1-1/5)=(3/5)/(4/5)=3/4